# Activation Steering in Neural Theorem Provers

### v1

@@ 4 Results and Analysis
Setup We evaluated the technique using Best First Search. It is one of the most popular methods to evaluate the theorem-proving ability of a language model Polu & Sutskever (2020b); Yang et al. (2023b); Azerbayev et al. (2023); Lin et al. (2024). For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$. Following standard practice Polu & Sutskever (2020b); Yang et al. (2023b); Welleck & Saha (2023); Lin et al. (2024) we assume the state with maximum negative log-probabilities is the ”best” state. Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average log probability of each generated token. We expand upto $N$ states and we get successful proofs search when we reach any proof state with no goals.
@@ 5 Conclusion
We investigate activation steering for tactic prediction by making language models adhere to structured reasoning approaches in theorem proving. We find that by constructing steering pairs using synthetic metadata and natural proof states, we can construct effective steering vectors that improve tactic selection. Our experiments show that steering vectors enhance model performance beyond random interventions and generalize well across different theorem-proving strategies. The effectiveness of our steering approach demonstrates the existence of underlying reasoning pathways that can be systematically influenced within language models. Activation steering proves to be a powerful technique for improving model performance on formal reasoning tasks where fine-tuning may be impractical or resource-intensive. As language models continue to evolve in their theorem-proving capabilities, activation steering may serve as a lightweight alternative to specialized fine-tuning approaches. Rather than training separate models for different theorem-proving strategies, the same base model could be adapted through steering vectors that guide its reasoning process as needed. This could be particularly valuable for interactive theorem proving environments where computational resources are limited. Our reasoning-based steering vectors, for example, could provide an efficient way to enhance proof assistants, particularly useful for applications like automated tactic suggestion. In future work, we aim to study the underlying mechanisms in language models responsible for structured mathematical reasoning. We further wish to explore how reasoning-focused steering vectors may generalize to open-ended theorem proving tasks and investigate their potential for improving proof search strategies.

### v3

@@ 4 Results and Analysis
Setup We evaluated the technique using Best First Search. It is one of the most popular methods to evaluate the theorem proving ability of a language model Polu & Sutskever (2020b); Yang et al. (2023b); Azerbayev et al. (2023); Lin et al. (2024). For a given language model $M$, we keep all unexpanded states $s_{i}$; each time we expand the best state $s_{i}$ and use the language model to sample $S$ net tactics $a_{i,1...S}$ for the current state $s_{i}$. Following standard practice Polu & Sutskever (2020b); Yang et al. (2023b); Welleck & Saha (2023); Lin et al. (2024) we assume the state with maximum negative log-probabilities is the ”best” state. Specifically, we select state $s_{i}$ with maximum $\sum_{j=0}^{i-1}-\log p(a_{j},s_{j})$, where $(s_{0},a_{0}),...,(s_{i-1},a_{i-1})$ is the proof trajectory before state $s_{i}$ and $\log p(a_{j},s_{j})$ is the average logarithmic probability of each generated token. We expand up to $N$ states and we get a successful proofs search when we reach any proof state with no goals.
  1. 1.
    patch scope
  2. 2.
    what sort of pathces are generated at each layer
  3. 3.
    what does random vector mean? explain how did we make it?
  4. 4.
    explain how changinng epsilon effect the generation in terms of effect of steering (should be a normal curve)
  5. 5.
    finetuning on P1 prompts,
  6. 6.
    benchmark with other models as well since these models are much smaller
  7. 7.
    activation steering with execution feedback like CORPA Thakur et al. (2024)
  8. 8.
    Proof Length and time to verify proofs.
  9. 9.
    add a dataset probably putnam bench and also ablations on how differnt proofs are from each other.
@@ 5 Conclusion
We investigate activation steering for tactic prediction by making language models adhere to structured reasoning approaches in theorem proving. We find that by constructing steering pairs using synthetic metadata and natural proof states, we can construct effective steering vectors that improve tactic selection. Our experiments show that steering vectors enhance model performance beyond random interventions and generalize well across different theorem-proving strategies. The effectiveness of our steering approach demonstrates the existence of underlying reasoning pathways that can be systematically influenced within language models. Activation steering proves to be a powerful technique for improving model performance on formal reasoning tasks where fine-tuning may be impractical or resource-intensive. As language models continue to evolve in their theorem-proving capabilities, activation steering may serve as a lightweight alternative to specialized fine-tuning approaches. This could be particularly valuable for interactive theorem proving environments where computational resources are limited. Our reasoning-based steering vectors, for example, could provide an efficient way to enhance proof assistants, particularly useful for applications like automated tactic suggestion. In future work, we aim to study the underlying mechanisms in language models responsible for structured mathematical reasoning. We further wish to explore how reasoning-focused steering vectors may generalize to open-ended theorem proving tasks and investigate their potential for improving proof search strategies.